OpenStudy (anonymous):

need help factoring: -20x *squared* + 29x - 6

OpenStudy (anonymous):

\[-20x^2+29x-6\]\[roots = \frac{(-b \pm \sqrt{b^2-4ac})}{(2a)} \] solve once with the + and other with - it will give you two values x1 and x2, use them for factor form: \[(x-x1)(x-x2)\]

OpenStudy (anonymous):

i still dont get it...

OpenStudy (anonymous):

\[ax^2+bx+c=0\]\[-20x^2+29x-6=0\] does it makes more sense now ?

OpenStudy (anonymous):

so how do you solve it?

hero (hero):

-(20x^2-29x+6) -(20x^2-24x-5x+6) -(-4x(-5x+6)+1(-5x+6)) -((-5x+6)(-4x+1))